OpenStudy (anonymous):

an atom has a mass number of 53 and an atomic number of 26. how many neutrons does it have? include a formula that explains your calculation

OpenStudy (anonymous):

The mass number of an atom, in general, is the amount of protons plus the amount of neutrons which make up its nucleus. The atomic number of an atom is the number of protons. So the number of neutrons, then, is equal to the mass number minus the atomic number, which in this case is 53-26=27 neutrons.

OpenStudy (whpalmer4):

Z = atomic number (number of protons, defines which element it is) A = mass number (number of protons and neutrons, also known as nucleons) N = neutron number (number of neutrons) A = Z+N N = A-Z